# Multiperiodic Processes: Ergodic Sources with a Sublinear Entropy

### v3

@@
Key words: ergodic processes, periodic sequences, zero entropy rate, Zipf‘s law, Hilberg‘s law
@@ 1 Introduction
The notations to state our problem are as follows. Blocks of random variables are denoted $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$. Let $H(X):=\operatorname{\mathbf{E}}\left[-\log P(X)\right]$ be the Shannon entropy of a discrete random variable $X$ where $\log$ is the base 2 logarithm, let H(X|Y):=H(X,Y)H(Y)H(X|Y):=H(X,Y)-H(Y) be the conditional entropy, and let I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y) be the mutual information [11]. Relation $\sim$ denotes is defined as
@@
Supplementing prior results, we seek for various toy examples of stationary processes $(X_{t})_{t\in\mathbb{Z}}$ that obey relationship
@@
where $\beta\in(0,1)$ is called the Hilberg exponent [19] and limit
$h:=\lim_{t\to\infty}\frac{H(X_{1}^{t})}{t}$ (3)
is the Shannon entropy rate. Relationship (2) is called Hilberg‘s law after [43]. The particular aim of the present paper is to construct simple examples of stationary ergodic (but not mixing) sources over a countable alphabet that obey Hilberg‘s law (2) with the vanishing entropy rate, $h=0$. The processes to be constructed are called multiperiodic processes and will be described after sketching the background.
To connect our problem to prior research, we note that limit
$E:=\lim_{t\to\infty}\left(H(X_{1}^{t})-ht\right)=\lim_{t\to\infty}I(X_{-t+1}^{0};X_{1}^{t})$ (4)
is called excess entropy [12] or predictive information [4, 3]. Excess entropy is an intuitive measure of memory stored in a stochastic process. Although this quantity only measures the memory capacity, without specifying how exactly the process future depends on the past, it can be given interesting interpretations. Some attempts were also made to generalize excess entropy to two-dimensional random fields [32, 8].
In fact, conditions $E<\infty$ and $E=\infty$ have been studied in probability theory and its applications for quite a time. The oldest systematic results seem to concern Gaussian processes. In particular, Grenander and Szegő [38, Section 5.5] present an integral formula for excess entropy (in disguise). Finch [33] evaluated this formula for autoregressive moving average (ARMA) processes, which turned out to yield $E<\infty$. More generally, under a positive and continuous spectral density, Li [49] showed that $E<\infty$ if and only if the autocorrelation $\rho_{k}:=\operatorname{Corr}(X_{0};X_{k})$ satisfies
$\sum_{k=1}^{\infty}k\rho_{k}^{2}<\infty.$ (5)
Dropping the condition for spectral density, we have $E<\infty$ if and only if $\sum_{k=1}^{\infty}k\alpha_{k}^{2}<\infty$ for the partial autocorrelation $\alpha_{k}$ [13].
Let us proceed to the countable alphabet. It can be easily shown, by the data-processing inequality, that condition $E<\infty$ holds for hidden Markov processes with a finite number of states [12]. Mahoney, Ellison and Crutchfield [52, 31] developed a formula for excess entropy in terms of predictive and retrodictive $\epsilon$-machines, which are minimal unifilar hidden Markov representations of the process [57, 51]. Travers and Crutchfield [58] and Dębowski [18] exhibited examples hidden Markov processes with a countably infinite number of hidden states that obey $E=\infty$ and (2). A few other older papers concerned also isolated examples of processes with $E=\infty$ over a finite alphabet. For instance, Bradley [7] constructed the first example of a mixing process having this property. Gramss [36] investigated a process which is formed by the frequencies of 0‘s and 1‘s in the rabbit sequence.
A more systematic study of condition $E=\infty$ for the countable alphabet case was inspired by the publication by Hilberg [43] suggesting that relationship (2) holds for natural language. As it will be detailed in a separate paper (to apply some results of the present one), Hilberg‘s law implies the more famous neural scaling law, which is a power-law decay of the cross entropy rate of large language models and empirically holds for internet-sized corpora of texts [47, 40, 42]. Although the first reception of Hilberg‘s hypothesis came from physicists [30, 29, 3, 12, 51], there is a line of research that seeks to combine a rigorous information-theoretic approach to Hilberg‘s law with general linguistic insights [14, 15, 16, 17, 24, 25] such as Zipf‘s law in particular [60, 53]. The basic observation is that condition $E=\infty$ holds if the process is strongly non-ergodic in the sense that its shift-invariant algebra $\mathcal{I}$ is non-atomic. It is so because
@@
The ergodic decomposition of excess entropy (6) gives rise to a family of linguistically motivated examples of processes that obey Hilberg‘s law (2). These examples are called Santa Fe processes [15, 16] and were independently considered also by Hutter [44]. We recall this family of examples at length for their certain structural similarities and differences to the multiperiodic processes constructed in this paper. The Santa Fe processes come in a number of varieties—some are strongly non-ergodic, some are ergodic. Their common simple idea is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. In other words, each text token $X_{t}$ may be written as a pair
@@
The Santa Fe decomposition (7) is interesting for its abstract semantic interpretation, discussed in [16, 24]. We may interpret that pairs $(K_{t},Z_{K_{t}})$ are logical statements that describe a certain unknown binary sequence $(Z_{k})_{k\in\mathbb{N}}$ at random but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may interpret that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable general knowledge which is only partially accessed and described in finite texts. [^1] There are potential connections between the Santa Fe decomposition (7) and algorithmic information theory [50], see also [24, Chapter 1]. In the role of immutable knowledge, the random probability $Y:=\sum_{k=1}^{\infty}2^{-k}Z_{k}\in(0,1)$ conceptually resembles Chaitin’s number $\Omega\in(0,1)$, also called the halting probability [9, 10]. The number $\Omega$ is algorithmically random [54] and encodes a large body of mathematical knowledge in a dense form. However, equating the numerical values of probabilities $Y$ and $\Omega$ is preposterous for two reasons. The first of them, a formal one, is that $Y$ is a random variable, whereas $\Omega$ is a fixed number. The second reason, philosophically a more important one, is that $\Omega$ does not seem computable in a sufficiently strong sense to arise as an effect of natural processes [2]. For this reason, bits $Z_{k}$ appearing in representation (7) were called the ’’(elementary) facts‘‘ [16], whereas sequence $(K_{t})_{t\in\mathbb{Z}}$ was called the ’’narration (process)‘‘ [24]. Thus, we may say that formula (7) decomposes a corpus of texts $(X_{t})_{t\in\mathbb{Z}}$ into ’’knowledge‘‘ $(Z_{k})_{k\in\mathbb{N}}$ and ’’narration‘‘ $(K_{t})_{t\in\mathbb{Z}}$. To complete the naming, individual variables $K_{t}$ might be called ’’topics‘‘.
A few previous discussions of Santa Fe processes assumed that narration $(K_{t})_{t\in\mathbb{Z}}$ is a memoryless source with the marginal Zipf-Mandelbrot distribution
@@
where $\alpha>1$ [60, 53]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (2) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^2] We adopt the distinction of terms “type” and “token” borrowed from quantitative linguistics [41, 48, 1]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
@@
For the Zipf-Mandelbrot distribution (8), we have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha}$ [41, 39, 1]. Consequently, we may express
$H(X_{1}^{t})-H(K_{1}^{t})=H(\left\{(k,Z_{k}):k\in\mathcal{V}_{t}\right\}|K_{1}^{t})=\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha},$ (10)
@@
To complete the collection of invented sources, we have been interested in finding a similarly simple example of a process with Hilberg‘s law (2) and the vanishing Shannon entropy rate, $h=0$. In [21, 24], we have constructed the random hierarchical association processes (RHA), which satisfy (2) with $h=0$ but are probably non-ergodic. The construction of RHA processes is so complicated that we do not reproduce it here. In the end, we desired a simpler construction that would be additionally ergodic.
Our curiosity in such asymptotically deterministic models was initially driven by the observation that the maximal repetition length for natural language obeys a cube logarithmic growth [20]. Consequently, the conditional Rényi entropy rate may vanish for natural language [22, 27]. Later, our interest was reinforced by the research in the neural scaling law [47, 40, 42], which suggested that Hilberg‘s law (2) holds for large language models with $h=0$. We are agnostic about the actual value of particular entropy rates for natural language but we have wished to better understand the mathematical construction of zero entropy rate processes.
@@
The essential idea of these processes is to define a probability measure on a set of efficiently computable but randomly shifted deterministic sequences, called multiperiodic sequences. Under a certain parameterization, the relative frequencies of particular numbers in these sequences approach the Zipf-Mandelbrot distribution (8). This suffices to imply Hilberg‘s law (11) with exponent $\beta=1/\alpha$ by a reasoning similar to bound (10). Namely, for the set of types (9), we still have the Herdan-Heaps law $\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha}$ and we may express
$H(K_{1}^{t})=H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\sim\operatorname{\mathbf{E}}\operatorname{card}\mathcal{V}_{t}\sim t^{1/\alpha},$ (12)
@@ 2.1 Multiperiodic sequences
Multiperiodic sequences are an idea generalized from the work by Kalociński and Steifer [46].
@@
Example 1 (seemingly A001511 in [55]).
For a solution of a different problem in prediction, Kalociński and Steifer [46] considered a particular multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ with constant periods $\pi_{k}=2$ and seeds $\sigma_{k}=1$, namely,
@@
As we show in Section 5.2, for periods $\pi_{k}\approx ck$, the relative frequency of number $k$ approaches the Zipf-Mandelbrot law (8) with $\alpha=(c+1)/c$.
@@ 2.2 Multiperiodic processes
Multiperiodic sequences can be applied to define stationary ergodic processes. The idea is to apply a random initialization, that is, uniformly random seeds. This leads to this construction, called a multiperiodic process.
@@
We have a simple criterion of existence of a multiperiodic process.
@@
Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [5, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
@@
Moreover, each multiperiodic process admits a frequency interpretation of its probability distribution. Namely, it is stationary and ergodic.
@@
for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [24, Section 4.1].
@@
Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [24, Theorem 4.10] or [37, Lemma 7.15]. ∎

### v4

@@
Key words: ergodic processes, periodic sequences, entropy rate, Zipf‘s law, Hilberg‘s law
@@ 1 Introduction
Supplementing prior instances such as Santa Fe processes introduced in [1, 2], the aim of the present paper is to construct further examples of stationary sources over a countable alphabet that obey a power-law growth of block entropy, called Hilberg‘s law [3]. The processes constructed in this paper are called multiperiodic processes. Multiperiodic processes are stationary ergodic (but not mixing) and enjoy the vanishing entropy rate—in contrast to Santa Fe processes [1, 2], which are non-ergodic and usually have a positive entropy rate. As it has been demonstrated in [4], Hilberg‘s law—deemed to hold for natural language [3]—implies the better known neural scaling law, which is a power-law decay of the cross entropy rate and empirically holds for large language models trained on internet-sized corpora of texts [5, 6, 7, 8]. We hope that the toy examples of Santa Fe and multiperiodic processes may be useful as baseline artificial data sources to be compared with natural texts in quantitative linguistics and machine learning.
We cast our contribution against the background of previous research.
Preliminaries.
The notations to state our problem are as follows. Blocks of random variables are denoted $X_{j}^{k}:=(X_{j},X_{j+1},\ldots,X_{k})$. Let $H(X):=\operatorname{\mathbf{E}}\left(-\log P(X)\right)$ be the Shannon entropy of a discrete random variable $X$ where $\log$ is the base 2 logarithm, let H(X|Y):=H(X,Y)H(Y)H(X|Y):=H(X,Y)-H(Y) be the conditional entropy, and let I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X|Y) be the mutual information. Relation $\sim$ is defined as
@@
For a stationary process $(X_{t})_{t\in\mathbb{Z}}$, we may define two limits
$h$ $:=\lim_{t\to\infty}\frac{H(X_{1}^{t})}{t}=\lim_{t\to\infty}H(X_{1}|X_{-t+1}^{0}),$ (2)
$E$ $:=\lim_{t\to\infty}\left(H(X_{1}^{t})-ht\right)=\lim_{t\to\infty}I(X_{-t+1}^{0};X_{1}^{t}).$ (3)
Limit $h$ is called entropy rate [9, 10], whereas $E$ is called excess entropy [11] or predictive information [12, 13, 14]. While entropy rate is a measure of process unpredictability, excess entropy is an intuitive measure of process memory. Although excess entropy only measures the memory capacity, without specifying how exactly the process future depends on the past, it can be given interesting interpretations. Some attempts were also made to generalize excess entropy to two-dimensional random fields [15, 16].
Memory conditions.
Conditions $E<\infty$ and $E=\infty$ have been studied in probability theory and its applications for quite a time. The oldest systematic results seem to concern Gaussian processes. In particular, Grenander and Szegő [17, Section 5.5] presented an integral formula for excess entropy (in disguise). Finch [18] evaluated this formula for autoregressive moving average (ARMA) processes, which turned out to yield $E<\infty$. More generally, we have $E<\infty$ if and only if $\sum_{k=1}^{\infty}k\alpha_{k}^{2}<\infty$ and $\left|\alpha_{k}\right|<1$ for the partial autocorrelation $\alpha_{k}$ [19]. In particular, under a positive and continuous spectral density, Li [20] showed that $E<\infty$ if and only if the autocorrelation $\rho_{k}:=\operatorname{Corr}(X_{0};X_{k})$ satisfies $\sum_{k=1}^{\infty}k\rho_{k}^{2}<\infty$.
Let us proceed to the countable alphabet. It can be easily shown, by the data-processing inequality, that condition $E<\infty$ holds for hidden Markov processes with a finite number of states [11]. Mahoney, Ellison and Crutchfield [21, 22] developed a formula for excess entropy in terms of predictive and retrodictive $\epsilon$-machines, which are minimal unifilar hidden Markov representations of the process [23, 24]. Travers and Crutchfield [25] and Dębowski [26] exhibited examples of hidden Markov processes with a countably infinite number of hidden states that obey $E=\infty$ and the further discussed Hilberg‘s law (4). A few other older papers concerned also isolated examples of processes with $E=\infty$ over a finite alphabet. For instance, Bradley [27] constructed the first example of a mixing process having this property. Gramss [28] investigated a process which is formed by the frequencies of 0‘s and 1‘s in the rabbit sequence.
A more systematic study of condition $E=\infty$ for the countable alphabet case was inspired by the publication by Hilberg [3] who redrew the plot by Shannon [29] in the log-log scale and proposed that relationship
@@
where $\beta\in(0,1)$, holds for natural language. Although the first reception of Hilberg‘s law (4) came from physicists [30, 31, 13, 11, 24], there is a line of research that seeks to combine a rigorous information-theoretic approach to Hilberg‘s law with general linguistic insights [32, 1, 2, 33, 34, 35] such as Zipf‘s law in particular [36, 37]. The basic observation is that condition $E=\infty$ holds if the process is strongly non-ergodic in the sense that its shift-invariant algebra $\mathcal{I}$ is non-atomic. It is so because
@@
Santa Fe processes.
The ergodic decomposition of excess entropy (5) inspires a family of linguistically motivated examples of processes that obey Hilberg‘s law (4). These examples are called Santa Fe processes [1, 2] and were independently considered by Hutter [43]. We recall this family of examples at length for their certain structural similarities and differences to the multiperiodic processes constructed in this paper. The Santa Fe processes come in a number of varieties—some are strongly non-ergodic, some are ergodic. Their common simple idea is to decompose each token $X_{t}$ as a pair of a natural number $K_{t}$ and an additional bit—which is copied from a certain sequence of bits $(Z_{k})_{k\in\mathbb{N}}$ by taking the bit at position $K_{t}$. In other words, each text token $X_{t}$ may be written as a pair
@@
The Santa Fe decomposition (6) is interesting for its semantic interpretation, discussed in [2, 34]. We may interpret that pairs $(K_{t},Z_{K_{t}})$ are logical statements that describe a certain binary sequence $(Z_{k})_{k\in\mathbb{N}}$ at random but in a non-contradictory way. Namely, if statements $(k,Z_{k})$ and $(k^{\prime},Z_{k^{\prime}})$ describe the same bit $(k=k^{\prime})$ then they assert the same value $(Z_{k}=Z_{k^{\prime}})$. Thus, we may interpret that sequence $(Z_{k})_{k\in\mathbb{N}}$ expresses some unbounded immutable knowledge which is only partially accessed and described in finite texts. For this reason, bits $Z_{k}$ appearing in representation (6) were called the ’’(elementary) facts‘‘ [2], whereas sequence $(K_{t})_{t\in\mathbb{Z}}$ was called the ’’narration (process)‘‘ [34]. Thus, we may say that formula (6) decomposes a corpus of texts $(X_{t})_{t\in\mathbb{Z}}$ into ’’knowledge‘‘ $(Z_{k})_{k\in\mathbb{N}}$ and ’’narration‘‘ $(K_{t})_{t\in\mathbb{Z}}$. To complete the naming, individual variables $K_{t}$ might be called ’’topics‘‘.
There may exist connections between the Santa Fe decomposition (6) and algorithmic information theory [44], see also [34, Chapter 1]. In the role of immutable knowledge, the random probability $Y:=\sum_{k=1}^{\infty}2^{-k}Z_{k}\in(0,1)$ resembles Chaitin‘s number $\Omega\in(0,1)$, also called the halting probability [45, 46]. The number $\Omega$ is algorithmically random [47] and encodes a large body of mathematical knowledge in a dense form. However, equating the values of probabilities $Y$ and $\Omega$ is preposterous for two reasons. The first of them, a formal one, is that $Y$ is a random variable, whereas $\Omega$ is a fixed number. The second reason, philosophically more important, is that $\Omega$ does not seem computable in a sufficiently strong sense to arise as an effect of natural processes [48]. If there are some links between the Santa Fe decomposition (6) and algorithmic information theory, they must be more subtle.
The default example of Santa Fe processes assumes that narration $(K_{t})_{t\in\mathbb{Z}}$ is a memoryless source with the marginal Zipf distribution
@@
where $\alpha>1$ [36, 37]. Moreover, it was assumed that knowledge $(Z_{k})_{k\in\mathbb{N}}$ is a sequence of independent fair coin flips, independent of process $(K_{t})_{t\in\mathbb{Z}}$. Under these conditions, we obtain Hilberg‘s law (4) with $\beta=1/\alpha$. The reason for this law is as follows. Denote the random set of types [^1] We adopt the distinction of terms “type” and “token” borrowed from quantitative linguistics [49, 50]: A token is an element of a sequence with possible repetitions, whereas a type is an element of a set, which excludes repetitions.
@@
For memoryless sources with Zipf‘s distribution (7), under mild conditions [51, 52], we have Heaps‘ law
$\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha}$ (9)
[49, 53, 50]. Consequently, we may express
$H(X_{1}^{t})-H(K_{1}^{t})=H(\left\{(k,Z_{k}):k\in\mathcal{V}_{t}\right\}|K_{1}^{t})=\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (10)
@@
Further examples.
To complete the collection of invented sources, we have been interested in finding an example of a process with Hilberg‘s law (4) and the vanishing Shannon entropy rate, $h=0$. In [54, 34], we have constructed the random hierarchical association processes (RHA), which satisfy (4) with $h=0$ but are probably non-ergodic. The construction of RHA processes is so complicated that we do not reproduce it here. In the end, we desired a simpler construction that would be also ergodic.
Our curiosity in such asymptotically deterministic models was initially driven by the observation that the maximal repetition length for natural language obeys a cube logarithmic growth [55]. Consequently, the conditional Rényi entropy rate may vanish for natural language [56, 57]. Later, our interest was reinforced by the research in the neural scaling law [5, 6, 7, 8], which suggested that Hilberg‘s law (4) holds for large language models with $h=0$. We are agnostic about the actual value of particular entropy rates for natural language but we have wished to better understand the mathematical construction of zero entropy rate processes.
Present contribution.
@@
The essential idea of these processes is to define a probability measure on a set of efficiently computable but randomly shifted deterministic sequences, called multiperiodic sequences. Under a certain parameterization, the relative frequencies of particular numbers in these sequences approach Zipf‘s distribution (7). This suffices to imply Hilberg‘s law (11) with exponent $\beta=1/\alpha$ by a reasoning similar to bound (10). Namely, for the set of types (8), we also have Heaps‘ law (9) and we may express
$H(K_{1}^{t})=H(\left\{(k,\Sigma_{k}):k\in\mathcal{V}_{t}\right\})\sim\operatorname{\mathbf{E}}\operatorname{\#}\mathcal{V}_{t}\sim t^{1/\alpha},$ (12)
@@ 2 Definitions
In this section, we construct two kinds of objects. The first one are multiperiodic sequences and the second one are muliperiodic processes. We also prove first theorems concerning these constructions: time reversal symmetry, periodicity of decimation, existence conditions, stationarity, and ergodicity. These properties will be used later for bounding some quantities of interest.
@@ 2.1 Multiperiodic sequences
Multiperiodic sequences are an idea which we have originally generalized from work [58]. The authors only considered a fixed multiperiodic sequence with particular periods and seeds—to state a certain negative result in universal prediction. Isolated instances of multiperiodic sequences are noted in the on-line encyclopedia of integer sequences [59] and can be connected to other problems, such as Mancala-type games [60]. Distilling the general construction, especially considering the idea of random seeds that matters in the probabilistic context, seems our contribution.
@@
Example 1 (A001511 in [59]).
Consider the multiperiodic sequence $(k_{t})_{t\in\mathbb{N}}$ with constant periods $\pi_{k}=2$ and seeds $\sigma_{k}=1$, namely,
@@
Equivalently, $k_{t}$ is the number of 22‘s dividing $2t$. The equivalence stems from the observation that $k_{t}$ is the position in the binary expansion of $t$ that is incremented when increasing from t1t-1 to $t$. This sequence was applied by Kalociński and Steifer [58] for a negative result in universal prediction.
@@
This sequence was discussed by Broline and Loeb [60] in the context of Mancala-type games. The algorithm for generating this sequence is isomorphic with recursion (13)–(17).
As we show in Section 5.2, for periods $\pi_{k}\approx ck$, the relative frequency of number $k$ approaches Zipf‘s law (7) with $\alpha=(c+1)/c$.
@@ 2.2 Multiperiodic processes
Multiperiodic sequences can be applied to define stationary ergodic processes. The basic idea is to apply a random initialization, that is, uniformly random seeds. This leads to the next construction, called a multiperiodic process. We observe that the existence of the process is conditional—given that the value $1$ appears in the seed sequence infinitely often almost surely.
@@
Applying the complementary Borel-Cantelli lemmas for independent events, we derive a simple criterion of existence of a multiperiodic process.
@@
Observe that events $(\Sigma_{k}=1)$ are independent. Hence by the Borel-Cantelli lemmas [61, Theorems 4.3 and 4.4], set $\left\{k\in\mathbb{N}:\Sigma_{k}=1\right\}$ is infinite almost surely if and only if
@@
Moreover, through the approximation by periodic processes, each multiperiodic process admits a frequency interpretation of its probability distribution. Namely, it inherits the properties of being stationary and ergodic from the approximating periodic sources.
@@
for a certain $p$-periodic sequence $(a_{t})_{t\in\mathbb{Z}}$. It is known that $p$-periodic processes are stationary and ergodic but they are not mixing [34, Section 4.1].
@@
Thus process $(K^{\leq m}_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. Since $m\in\mathbb{N}$ is arbitrarily chosen, hence also process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and ergodic but is not mixing. The reason is that verifying stationarity, ergodicity, and mixing can be reduced to checking probabilities of finite blocks by [34, Theorem 4.10] or [62, Lemma 7.15]. ∎